OpenStudy (anonymous):

Find the center of the circle whose equation is x² - 10x + y² - 10y = -1.

OpenStudy (anonymous):

x^2-10x+25-25+y^2-10y+25-25=-1 (x-5)^2+(y-5)^2-25-25=-1 (x-5)^2+(y-5)^2=49 you know how to find center of circle from this equation?